Consider the demand equation below, where x represents the quantity demanded in units of a thousand and p is the unit price in dollars.
p = -x2 + 11
(a) Determine the quantity demanded when the unit price is set at p = 7.

units

Found 2 solutions by rfer, stanbon:
Answer by rfer(16322)

7=-x^2+11
7-11=-x^2
-4=-x^2
x=2

Answer by stanbon(75887)

Consider the demand equation below, where x represents the quantity demanded in units of a thousand and p is the unit price in dollars.
p = -x2 + 11
(a) Determine the quantity demanded when the unit price is set at p = 7.
------
Solve for "x":
x^2+p = 11
x^2 = 11-p
x = sqrt(11-p)
----
If p = 7
x = sqrt(11-7)
x = sqrt(4)
x = 2
=============
Cheers,
Stan H.
==============